Chrissie's Cooking Supply Company has 5,000 skillets in their warehouse at the end of July. One-quarter of these skillets were held over from the month of June at a cost of $12 per skillet. The remaining skillets were purchased in July at a cost of $15 per skillet. At the beginning of August, they received another 2,000 skillets at a cost of $17 per skillet. The warehouse sold and shipped 2,198 skillets during August. Chrissie's Cooking Supply Company uses LIFO to value their inventory. What would be the remaining balance of skillets in the inventory account at the end of August?

Explain answer

1.$68,280

2.$76,030

3.$29,220

4.$36,970
Answer #1

Beginning inventory = (5000*1/4*12+5000*3/4*15) = 71250

Purchase = 2000*17 = 34000

Cost of goods sold = (2000*17+198*15) = 36970

Ending inventory = 71250+34000-36970 = $68280

So answer is a) $68280